OpenStudy (anonymous):

x/a+x/b=c; where a is not equal to 0, b is not equal to zero, a is not equal to -b, solve for x.

OpenStudy (anonymous):

\[\frac{x}{a}+\frac{x}{b}=c\] \[a \neq 0, b \neq 0, a \neq -b\]

OpenStudy (anonymous):

it doesn't say :(

OpenStudy (anonymous):

what i listed is all that was given /:

OpenStudy (anonymous):

\[x \left( \frac{ 1 }{ a } +\frac{ 1 }{ b }\right)=c\] \[x \left( \frac{ b+a }{ ab } \right)=c,x=?\]

OpenStudy (anonymous):

do i multiply x*b+a and x*ab??

OpenStudy (anonymous):

\[x=\frac{ cab }{ b+a }\]

OpenStudy (anonymous):

@surjithayer I'm sorry I'm a little confused.. is that the answer that you're giving or do i need to do something else from there? :o

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

to get the value of x,multiply both sides by ab and then divide both sides by b+a